OpenStudy (anonymous):

3. C is the midpoint of segment AB, and AC= 3x+7; BC= 5x - 13 solve for x @Mathematics

OpenStudy (anonymous):

These have to equal each other, 3x+7=5x-13

OpenStudy (anonymous):

is the answer 6.6 repeating?

OpenStudy (anonymous):

No. Double check your algebra. 3x - 5x = -13 - 7

OpenStudy (anonymous):

oooo i wrote it wrong

OpenStudy (anonymous):

10?